A crate with a mass of 22.0 kg is moved up a flight of stairs onto a landing. If the height of the landing is 3.00 meters, how much work was done?
1
Expert's answer
2020-11-05T10:12:05-0500

By the definition of work done, we have:


"W=Fh=mgh=22.0\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot 3.0\ m=647\ J."

Answer:

"W=647\ J."